A car travelling at 15 m/s comes to rest in a distance of 14 m when the brakes are applied.

Calculate the deceleration of the car.


1
Expert's answer
2021-07-21T09:04:44-0400

According to the work-energy theorem, the amount of work required to stop the car is equal to its initial kinetic energy:


"W = \dfrac{mv^2}{2}"

where "v = 15m\/s" and "m" is the mass of the car.

On the other hand, the work is given as follows:


"W = Fd"

where "d = 14m" and "F" is the friction force that were acting on the car. By the second Newton's law:


"F = ma"

where "a" is the decceleration of the car. Combining this all together, obtain:


"mad = \dfrac{mv^2}{2}\\
a = \dfrac{v^2}{2d}"

Thus, obtain:


"a = \dfrac{15^2}{2\cdot 14} \approx 0.57m/s^2"

Answer. 0.57 m/s^2.